# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

### v2

@@ Abstract
Aligning human preference and value is an important requirement for building contemporary foundation models and embodied AI. However, popular approaches such as reinforcement learning with human feedback (RLHF) break down the task into successive stages, such as supervised fine-tuning (SFT), reward modeling (RM), and reinforcement learning (RL), each performing one specific learning task. Such a sequential approach results in serious issues such as significant under-utilization of data and distribution mismatch between the learned reward model and generated policy, which eventually lead to poor alignment performance. We develop a single stage approach named Alignment with Integrated Human Feedback (AIHF), capable of integrating both human preference and demonstration to train reward models and the policy. The proposed approach admits a suite of efficient algorithms, which can easily reduce to, and leverage, popular alignment algorithms such as RLHF and Directly Policy Optimization (DPO), and only requires minor changes to the existing alignment pipelines. We demonstrate the efficiency of the proposed solutions with extensive experiments involving alignment problems in LLMs and robotic control problems in MuJoCo. We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF and DPO by large margins, especially when the amount of high-quality preference data is relatively limited.
[…]
@@ 1 Introduction
Background. As ChatGPT has taken the world by storm, it is clear that AI systems will soon become ubiquitous in our lives. For instance, Large Language Models (LLMs) have been used to solve hard problems including video gaming (Berner et al., 2019; Mnih et al., 2015), autonomous control (Bellemare et al., 2020), and robotic manipulation (Kalashnikov et al., 2018; Kober and Peters, 2008). However, for AI systems to be able to smoothly interact with humans, and reliably help perform mission-critical tasks, it is critical that the system learns, or aligns with, the human’s preferred strategy, under specific contexts, and adapts to evolving user preferences and needs (Leike et al., 2018). Achieving this level of alignment between AI and humans requires a comprehensive approach. AI systems must be capable of not only responding to instructions but also being sensitive to the nuances of human behavior and context, ensuring that their actions align with individual values and societal norms. On the other hand, various undesirable behaviors exhibited by advanced AI systems, e.g. deception (Park et al., 2023) and manipulation (Perez et al., 2022) have raised concerns about the potential hazards from AI systems.
The “alignment” task mentioned above has already proven to be useful in a few applications, such as in ChatGPT. Despite its recent popularity, the research area is still in its infancy. There are a number of critical open challenges, e.g. 1) How to solve the alignment problem leveraging diverse human feedback to further boost alignment performance? 2) How to enable AI systems to continuously learn and adapt to time-varying preferences, changing contexts, and emerging societal norms?
Contribution. In this work, we address the first question posed above. That is, we develop formulations and efficient algorithms capable of integrating different sources of human response data to achieve the best alignment performance.
@@
Most of the existing alignment algorithms are sequential, i.e. they separate the alignment process into a number of successive stages, each utilizing one distinctive category of data. For example, the well-known RLHF approach adopted by (Ouyang et al., 2022) breaks the alignment step into the supervised finetuning (SFT), reward modeling (RM), and reinforcement learning (RL) stages, where the SFT stage uses demonstration data, the RM stage uses the preference data, and the RL stage uses a set of available prompts. Similar strategies have been used in other related works such as (Rafailov et al., 2023; Li et al., 2023; Zhu et al., 2023; Liu et al., 2023). Sequencing of learning tasks may facilitate computation at the expense of inefficient exploitation of data. To illustrate, consider the three-stage RLHF approach proposed in (Ouyang et al., 2022), in the extreme case where the amount of high-quality preference data is quite limited. In such a situation, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage. This simple example underscores the importance of ensuring the reward model is consistent with the policy model. Such consistency can not be guaranteed with sequential execution of the RM and RL tasks.
The alignment problem is a learning problem with (at least) three types of input data: the demonstration data (consists of prompts and human-generated continuations), the preference data (consists of prompts and pairs of human-ranked responses), as well as prompts without any responses. The input to the learning problem, besides these data, also includes a model pre-trained using a large corpus of text data. The output of the learning process is a fine-tuned model aligned with human preferences/values which govern the generation of demonstration and preference data.
@@
To integrate these diverse data, we propose a joint learning approach, which is informed by recent advances in Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Zeng et al., 2022b), stochastic choice theory (Blavatskyy and Pogrebna, 2010) and bi-level optimization (Hong et al., 2020; Ji et al., 2021; Khanduri et al., 2021). More specifically, the proposed formulation integrates SFT, RM, and RL into a single stage, so that reward modeling and policy optimization can fully leverage all three kinds of data mentioned above in alignment. Several existing alignment schemes, such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023), and some of their extensions can be seen as particular instances of the proposed formulation. We observe that the proposed solutions outperform the existing alignment algorithms such as RLHF by large margins, especially when the amount of high-quality preference data is relatively limited.
@@ 2.1 Notation
The Finite-Horizon MDP Model. Consider a Markov decision process (MDP) defined by the tuple $(\mathcal{S},\mathcal{A},P,\rho,r,\gamma)$, which consists of the state space $\mathcal{S}$, the action space $\mathcal{A}$, the transition dynamics $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$, the initial state distribution $\rho(\cdot)$, the reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ and the discounted factor $\gamma>0$. Let $s_{t}\in\mathcal{S}$, $a_{t}\in\mathcal{A}$ denote the state and action variables at time $t$ respectively, where $a_{t}$ is generated according to some policy $\pi(\cdot|s_{t}):\mathcal{S}\to\Delta_{|\mathcal{A}|}$, where $\Delta_{|\mathcal{A}|}$ denotes the probability simplex in $\mathbb{R}^{|A|}$. Note that viewing the generation process of language models as an MDP, $s_{t}$’s and $a_{t}$’s are referred to as the given sequence and next token to be generated, respectively; $\pi(\cdot)$ then represents the language model itself. Define $\tau:=\{(s_{t},a_{t})\}_{t=1}^{T}$ as a (finite time) trajectory of state and action pairs. Let $\mathcal{H}_{T}\subset\prod_{t=0}^{T-1}\big{(}S\times A\big{)}$ denote all feasible state/action pairs.
Human Feedback Generation. Assume that the human demonstration $\tau$ is generated from a ground-truth human expert policy $\pi^{\rm E}$, i.e., for each state $s_{t}$, we have $a_{t}\sim\pi^{\rm E}(\cdot|s_{t})$. Moreover, let ‘$\tau_{l}\prec\tau_{w}$’ indicate the winner trajectory $\tau_{w}$ is preferred over the trajectory $\tau_{l}$. Further, we denote $(\tau_{l}\prec\tau_{w})\sim\pi^{\rm P}$ as the preference oracle. Note that when we present our problem formulation, we will use the above data generation distributions so quantities such as likelihood can be properly defined. Finite-sample version of demonstration and preference data will be introduced subsequently to facilitate algorithm design.
Since reward modeling is a critical task for achieving state-of-the-art alignment performance, below we review a few popular paradigms for reward learning.
2.2 Reward Learning from Preference Data
@@
RLHF is a popular technique for finetuning AI systems to align with human preferences and values. The RLHF approach proposed in (Stiennon et al., 2020; Ouyang et al., 2022) consists of the following three-stage: 1) the supervised fine-tuning (SFT) stage, where the demonstration data is used to fine-tune the model in a supervised manner; 2) the reward modeling (RM) stage, where the preference data is used to train a reward model; 3) the reinforcement learning (RL) stage, where the SFT model is further finetuned by running RL using the trained reward model.
Furthermore, existing work such as Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna and Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences. This greatly reduced the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024). There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022). However, these works often fail when the quality of the preference dataset is low Morimura et al. (2024). In our approach, the inclusion of additional demonstrations can be leveraged to counteract the adverse effects of limited preference quality.
@@ 2.3 Reward Learning using Demonstration Data
In the RL literature, a line of work referred to as Inverse Reinforcement Learning (IRL) proposes to jointly learn the reward and policy from expert demonstration data. Specifically, the target is to find the parameterized reward function $r(s,a;\theta)$ (resp. an optimal polity $\pi^{*}(s,a)$) that best explains (resp. mimics) an expert policy $\pi^{E}$ given the demonstration data $\mathcal{D}$. For example, the well-known maximum entropy IRL (MaxEnt-IRL) framework (Ziebart, 2010; Ziebart et al., 2013; Bloem and Bambos, 2014; Zhou et al., 2017) finds a policy maximizing entropy subject to the expected features that match the empirical averages in the expert’s observation dataset. However, this approach can only be used to model linear rewards.
@@
Subsequent works such as (Levine et al., 2011; Wulfmeier et al., 2015; Zeng et al., 2022b) further improve the MaxEnt-IRL method so that nonlinear reward can be used. For example, a recent paper (Zeng et al., 2022a) proposed a maximum likelihood IRL (ML-IRL) formulation based on the Dynamic Discrete Choice (DDC) model, and nonlinear reward function can be used. It is also shown that when the reward function is linearly parameterized, then the MaxEnt is the Lagrangian dual of the ML-IRL problem. However, it is worth mentioning that to our best knowledge, almost all IRL-based methods can only utilize demonstration data, which can be problematic because a large amount of high-quality demonstration data is typically hard to obtain. Further, it is well-known that only using demonstration cannot extract precise human preference, especially in safety-related tasks where the boundaries between permissible and impermissible actions need to be precisely determined (Liao and Sun, 2024); see, e.g., (Fischer et al., 2021) which shows that insufficient demonstration dataset could lead to high generalization error.
2.4 Reward Learning from Diverse Data
@@
For a more precise alignment of agent behavior with human intention, it is crucial to leverage diverse data source such as both expert demonstrations and human feedback to achieve alignment. In (Ibarz et al., 2018), the authors first combine two approaches to learn from human feedback: expert demonstrations and trajectory preferences. The addition of demonstrations to learning from preferences typically results in substantial performance gains compared with using either demonstrations or preferences in isolation. In (Palan et al., 2019) and (Bıyık et al., 2022), the authors integrate diverse sources of human feedback including demonstrations and pairwise comparisons in a Bayesian approach to learn reward functions that are assumed to be linear in carefully selected features and evaluate their proposed method on robot learning platform. Moreover, their proposed methods need to actively generate preference queries, which are expensive to collect in practical applications. In contrast, our proposed approach (to be presented shortly) is not Bayesian and does not include the requirement that the reward model is linear in pre-selected features. Hong et al. (2024) proposed a single-stage supervised learning algorithm ORPO that can perform supervised fine-tuning and preference alignment in one training session without maintaining multiple models simultaneously. However, Hong et al. (2024) can suffer from overoptimization due to its supervised learning pipeline Liu et al. (2024).
3 Proposed Approach: Alignment with Integrated Human Feedback (AIHF)

### v3

# Learning Reward and Policy Jointly from Demonstration and Preference Improves Alignment

@@ Abstract
Aligning to human preferences and/or intentions is an important requirement for contemporary foundation models. To ensure alignment, popular approaches such as reinforcement learning with human feedback (RLHF) break down the task into three stages: (i) a model is computed with supervised fine-tuning (SFT) based upon large demonstrations data, (ii) a reward model (RM) is estimated based upon human feedback data, and (iii) reinforcement learning (RL) is used to further refine the SFT model by optimizing the estimated reward model. Demonstrations and human feedback data reflect human user preferences in different ways. As a result, the reward model estimate obtained from only human feedback data is likely not as accurate as a reward model estimate obtained from both demonstration and human feedback data. A policy model that optimizes the reward model estimate obtained from both demonstration and human feedback data will likely exhibit better alignment performance. We introduce a tractable algorithm for finding the reward and policy models and provide a finite-time performance guarantee. Additionally, we demonstrate the efficiency of the proposed solution with extensive experiments including alignment problems in LLMs and robotic control problems in MuJoCo. We observe that the proposed solutions outperform the existing alignment algorithm by large margins.
@@ 1 Introduction
As ChatGPT has taken the world by storm, it is clear that AI systems will soon become ubiquitous in our lives. For instance, Large Language Models (LLMs) have been used to solve hard problems including video gaming (Berner et al., 2019; Mnih et al., 2015), autonomous control (Bellemare et al., 2020), and robotic manipulation (Kalashnikov et al., 2018; Kober & Peters, 2008). In this context, the notion of alignment plays an increasingly important role in the design and training of AI systems. Loosely speaking, alignment refers to the performance guarantee that the AI system will generate outcomes that are intended or preferred by the human user without undesirable side effects or behaviors such as deception (Park et al., 2023) or manipulation (Perez et al., 2022). As human user intentions or preferences may vary under specific contexts, it is critical that the AI system adapts to evolving user preferences and/or intentions (Leike et al., 2018).
@@
The alignment problem is a learning problem with (at least) three types of input data: the demonstration data (consists of prompts and human-generated continuations), the preference data (consists of prompts and pairs of human-ranked responses), as well as prompts without any responses. Moreover, the process of aligning an LLM model is typically undertaken in successive stages. For example, the well-known RLHF approach adopted by Ouyang et al. (2022) starts with a supervised fine-tuning model (SFT) followed by reward model (RM) estimation based upon human-labeled preference data. The process closes with a final alignment stage in which reinforcement learning (RL) is used to optimize the estimated reward model. Similar strategies have been used in other related works such as Rafailov et al. (2023); Li et al. (2023); Zhu et al. (2023); Liu et al. (2023). The approach to alignment based on successive stages may facilitate computation, but it is at the expense of inefficient exploitation of data. To illustrate, consider the three-stage RLHF approach proposed in Ouyang et al. (2022), in the extreme case where the amount of high-quality preference data is quite limited, the reward model trained cannot adequately reflect the preferences of the human, which may lead to unsatisfactory performance in the RL stage. Further, the reward model estimate obtained from only the preference data fails to exploit the information about human users’ preferences that are implicit in demonstration data. It is therefore reasonable to expect that a policy model that is fine-tuned with the reward model estimate obtained from both demonstration and human feedback may exhibit better alignment performance.
An alternative to the successive approach to alignment consists of jointly training the reward and policy models by leveraging demonstration and preference data. In contrast to the successive approach adopted in most of the current alignment approaches, the joint approach to reward and policy learning makes use of all available data, hence mitigating the risk of optimizing an inaccurate reward model. However, a joint approach to learning reward and policy models may improve alignment at the expense of potentially significant additional computational effort.
@@
Contribution. We introduce an algorithm jointly learning reward and policy models named Alignment with Integrated Human Feedback (AIHF) with a finite-time performance guarantee. This approach leverages recent advances in Inverse Reinforcement Learning (IRL) (Arora & Doshi, 2021; Zeng et al., 2022b), stochastic choice theory (Blavatskyy & Pogrebna, 2010) and bi-level optimization (Hong et al., 2020; Ji et al., 2021; Khanduri et al., 2021). The proposed formulation integrates SFT, RM, and RL into a single stage, so that reward modeling and policy optimization can fully leverage all the available human feedback data. More specifically, in the proposed algorithm, the policy is updated to improve alignment with the current reward model estimate and the reward model is updated to improve the fit to demonstration and human feedback data. As a result, upon convergence, the resulting reward and policy models are consistent in the sense that (i) the policy model is optimal with respect to the reward model and (ii) the reward model maximizes the fit to both demonstration and human feedback data. Several existing alignment schemes, such as RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023) and some of their extensions can be seen as particular instances of the proposed formulation. We provide ample empirical evidence that the proposed AIHF solution outperforms the existing alignment algorithms by large margins, especially when the data is unbalanced, where the quality and/or quantity of one data category is worse/smaller than that of the other.
@@ 2.1 Notation
The Finite-Horizon MDP Model. A Markov decision process (MDP) is the tuple $(\mathcal{S},\mathcal{A},P,\rho,r,\gamma)$, wherein $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the transition probabilities, $\rho(\cdot)$ is the initial state distribution, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function and $\gamma\in(0,1)$ denotes the discount factor. For every $s_{t}\in\mathcal{S}$, a randomized policy $\pi(\cdot|s_{t})$ is a probability distribution in $\Delta_{|\mathcal{A}|}$, the unit simplex in $\mathbb{R}^{|A|}$. Define $\tau:=\{(s_{t},a_{t})\}_{t=1}^{T}$ as a (finite horizon $T$) trajectory of state and action pairs. Let $\mathcal{H}_{T}\subset\prod_{t=1}^{T}\big{(}\mathcal{S}\times\mathcal{A}\big{)}$ denote all feasible state/action sequence of length $T$.
MDP Model of LLM. The generation of text by a language model can be seen as sampling from policies in an MDP model. Specifically, each state $s_{t}=(x,y_{1:t-1})$ includes the prompt $x$ and all response tokens produced up to that point $y_{1:t-1}$. Each action $a_{t}=y_{t}$ represents a token from the vocabulary. The transition kernel $P$ is deterministic, i.e. given tokens $s_{t}=(x,y_{1:t-1})$ and $a_{t}=y_{t}$, the environment will transition to $s_{t+1}=(x,y_{1:t})$. An LLM can be seen as a policy $\pi(\cdot|s_{t})$ so that a response of length $T>0$ to prompt $x$ is obtained with probability:$\pi(y_{1:T}|x):=\prod_{i=1}^{T}\pi(y_{i}|x,y_{1:i-1})$
Human Feedback Data. Let $\tau:=(y_{1:T},x)$ denote a finite text produced in response to prompt $x$. For a pair of sequences $(\tau_{l},\tau_{w})$ (which we assume of the same length $T$ for ease of exposition) we write $\tau_{l}\prec\tau_{w}$ to indicate the sequence $\tau_{w}$ is preferred over the sequence $\tau_{l}$. Following the Bradley-Terry-Luce (BTL) model (Bradley & Terry, 1952), the distribution of preferences over pairs $(\tau_{l},\tau_{w})$ can be modeled as follows:
$P\big{(}\tau_{w}\succ\tau_{l}\big{)}=\frac{\exp R(\tau_{w};\theta)}{\exp R(\tau_{w};\theta)+\exp R(\tau_{l};\theta)}=\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}$ (1)
where $\sigma$ is the sigmoid function and $R(\tau;\theta):=\sum_{t\geq 1}^{T}\gamma^{t}r(s_{t},a_{t};\theta)$ and $r(s_{t},a_{t};\theta)$ is a reward model parametrized by $\theta\in\mathbb{R}^{d}$.
2.2 The RLHF Pipeline
@@
RLHF is a popular technique for fine-tuning AI systems to align with human preferences and values. The RLHF approach proposed in Stiennon et al. (2020); Ouyang et al. (2022) consists of the following three-stage: 1) the supervised fine-tuning (SFT) stage, where the demonstration data is used to fine-tune the model in a supervised manner; 2) the reward modeling (RM) stage, where the preference data is used to train a reward model; 3) the reinforcement learning (RL) stage, where the SFT model is further fine-tuned by running RL using the trained reward model. Specifically, the RLHF pipeline can be formally described as follows:
Supervised Fine-Tuning (SFT): Given a demonstration dataset $\mathcal{D}$ consisting of sequences of the form $\tau=\{(s_{t},a_{t})\}_{t\geq 0}$ the goal is the find the policy $\pi_{\rm SFT}(\cdot|s_{t})$ that maximizes likelihood, i.e.:
$\pi_{\rm SFT}$ $=\arg\max_{\pi}\mathbb{E}_{\tau\sim\mathcal{D}}\Big{[}\log\prod_{t\geq 0}\Big{(}\pi(a_{t}|s_{t})\Big{)}^{\gamma^{t}}\Big{]}$ (2)
Reward Modeling (RM): Based upon a dataset $\mathcal{P}$ of preferences over pairs $(\tau_{l},\tau_{w})$ the estimation of a reward learning problem can be formulated as the following Bradley-Terry-Luce (BTL) model (Bradley & Terry, 1952) (with $\beta>0$ a hyper-parameter):
$\max_{\theta\in\mathbb{R}^{d}}\ell_{\rm RM}(\theta):=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\in\mathcal{P}}\Big{[}\log\Big{(}\sigma\big{(}\frac{1}{\beta}\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (3)
Reinforcement Learning (RL): Let $\hat{\theta}_{\mathcal{P}}$ denote the solution to problem (3). The last stage in the RLHF development pipeline consists of solving the problem:
$\pi_{\rm RLHF}$ $=\arg\max_{\pi}\mathbb{E}_{\tau\sim\pi}\Big{[}\sum_{t\geq 0}\gamma^{t}\big{[}r(s_{t},a_{t};\hat{\theta}_{\mathcal{P}})-\beta{D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}\Big{]}$ (4)
where ${D}_{\rm KL}\Big{(}\pi(\cdot|s_{t})\|\pi_{\rm SFT}(\cdot|s_{t})\Big{)}:=\sum_{a\in\mathcal{A}}\pi(a|s_{t})\log\frac{\pi(a|s_{t})}{\pi_{\rm SFT}(a|s_{t})}$ is the Kullback-Leibler (KL) divergence, $\pi_{\rm SFT}$ is the supervised fine-tuning model.
@@ 2.3 Reward Learning using Demonstration Data
In the RL literature, a line of work referred to as Inverse Reinforcement Learning (IRL) proposes to jointly learn the reward and policy from expert demonstration data. Specifically, the target is to find the parameterized reward function $r(s,a;\theta)$ (resp. an optimal polity $\pi^{*}(s,a)$) that best explains (resp. mimics) an expert policy $\pi^{E}$ given the demonstration data $\mathcal{D}$. For example, the well-known maximum entropy IRL (MaxEnt-IRL) framework (Ziebart, 2010; Ziebart et al., 2013; Bloem & Bambos, 2014; Zhou et al., 2017) finds a policy maximizing entropy subject to the expected features that match the empirical averages in the expert’s observation dataset. However, this approach can only be used to model linear rewards.
Subsequent works such as Levine et al. (2011); Wulfmeier et al. (2015); Zeng et al. (2022b) further improve the MaxEnt-IRL method so that nonlinear reward can be used. For example, Zeng et al. (2022a) proposed a maximum likelihood IRL (ML-IRL) formulation based on the Dynamic Discrete Choice (DDC) model, and a nonlinear reward function is utilized.
@@
It is also shown that when the reward function is linearly parameterized, then the MaxEnt is the Lagrangian dual of the ML-IRL problem (Zeng et al., 2022a). On the other hand, it is worth mentioning that to our best knowledge, almost all IRL-based methods can only utilize demonstration data, which can be problematic because a large amount of high-quality demonstration data is typically hard to obtain. Further, it is well-known that using demonstration only cannot extract precise human preference, especially in safety-related tasks where the boundaries between permissible and impermissible actions need to be precisely determined; see, e.g., Fischer et al. (2021) which shows that insufficient demonstration dataset could lead to high generalization error.
2.4 Joint Learning from demonstration and preference
@@
Combining data from demonstrations and human feedback to achieve alignment has also been studied in the robotics literature. In Ibarz et al. (2018), the authors first combine two approaches to learn from human feedback: expert demonstrations and trajectory preferences. The addition of demonstrations to learning from preferences typically results in substantial performance gains compared with using either demonstrations or preferences in isolation. In Palan et al. (2019) and Bıyık et al. (2022), the authors integrate diverse sources of human feedback including demonstrations and pairwise comparisons in a Bayesian approach to learn reward functions that are assumed to be linear in carefully selected features and evaluate their proposed method on robot learning platform. Moreover, their proposed methods need to actively generate preference queries, which are expensive to collect in practical applications. In contrast, the approach proposed in this paper is not Bayesian and does not include the requirement that the reward model is linear in pre-selected features.
2.5 Other Approaches to Alignment
Other approaches to alignment include Direct Preference optimization (DPO) (Rafailov et al., 2023) and Inverse Preference Learning (IPL) (Hejna & Sadigh, 2023) both remove the need for explicit reward modeling, and they directly extract the policy from preferences. This greatly reduced the training complexity, but it has been observed that these algorithms can be unstable in the training process (Azar et al., 2023; Xu et al., 2024). There is also a large number of works that aim to learn reward functions from rating (Daniel et al., 2014) or ranking (Yuan et al., 2023; Myers et al., 2022). Hong et al. (2024) proposed a single-stage supervised learning algorithm ORPO that can perform supervised fine-tuning and preference alignment in one training session without maintaining. However, all of these works highly rely on high-quality human feedback, which is often more difficult and expensive to obtain.
3 Alignment with Integrated Human Feedback (AIHF)
As mentioned before, the reward model obtained in (3) fails to exploit the information about human users’ preferences that are implicit in demonstration data. As a result, the fine-tuned model obtained with RLHF may exhibit unsatisfactory alignment performance (this phenomenon will be discussed more concretely in Sec. 3.4). Below we introduce a new approach to jointly train reward and policy models by simultaneously leveraging demonstration and human feedback data.